find the distance between
(-3,2) and (1,4)
Thank you

Answer by alka(15)

distance = sqrt[(1+3)^2 + (4-2)^2 ]
= sqrt(16 + 4)
= sqrt(20)
= sqrt(5*4)
= 2sqrt(5)